A company is making water bottles. The set-up cost is $200.00 and the cost per unit is $5.00. Each unit of the product will sell for $13.00. How many water bottles must they make to break even?
Answer by checkley77(12569)

200+5X=13X
200=13X-5X
8X=200
X=200/8
X=25 BOTTLES WILL BE A BREAK EVEN POINT.
PROOF:
200+5*25=13*25
200+125=325
325=325